OpenStudy (anonymous):

A chain of lenght 1m lying lenghtwise on a fixed table is partially hanging vertically outside the edge of the table . If the coefficient of friction b/w chain and the table is 0.5 the max lenght of the chain that can hang from the table without slipping?

OpenStudy (ghazi):

 let the mass per unit length of the chain = m then mass of length x = m*x and it's weight= m*x*g

OpenStudy (ghazi):

to find the maximum length...just equate weight of hanging chain= force of friction...because if is the force of friction that will stop chain from slipping and force of friction= u* normal= 0.5* m(1-x)*g

OpenStudy (ghazi):

your final equation is \[0.5*m(1-x)g= m*x*g\] hope this helps you :)

OpenStudy (ghazi):

@Yahoo! did you get 0.33m

OpenStudy (anonymous):

wat would be its mass

OpenStudy (ghazi):

you can't calculate mass ...don't worry it will be canceled on both sides

OpenStudy (ghazi):

@Yahoo! check out the above equation

OpenStudy (anonymous):

oh ...yup.....u explained this....so...well..lol

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

thxxx a lot

OpenStudy (ghazi):

you're welcome